OpenStudy (anonymous):

A set of data items is normally distributed with a mean of 60 and a standard deviation of 9. Convert 60 to a z-score. Round your answer to the nearest hundredth as needed.

OpenStudy (anonymous):

z=(x-(mean))/(standard deviation) here z= (60-60)/9 =0